OpenStudy (anonymous):

integrate dx/[(sqrt(x))(3^(sqrt(x))]

OpenStudy (anonymous):

I put u=sqrt(x)

hartnn (hartnn):

put u= sqrt x du=.. ?

OpenStudy (anonymous):

2du=1/sqrt(x)dx

OpenStudy (anonymous):

think i'm stuck where i need to use a log property..

hartnn (hartnn):

oh, you don't need to, don't you know the standard formula for \(\int a^xdx=... ?\)

OpenStudy (anonymous):

wouldn't I have 3^(-u) at one point though?

hartnn (hartnn):

yes, so you can use the formula for integral of a^x dx, right ?

hartnn (hartnn):

\(2\int 3^{-u}du = 2[\dfrac{3^{-u}}{(-1)\log u}]+c \)

OpenStudy (anonymous):

we can also do it by substituting 3^ sqrt(x) as t

OpenStudy (anonymous):

where did the log come from? what am i forgetting?

hartnn (hartnn):

\(\int a^xdx=a^x/ \log a +c\)

OpenStudy (anonymous):

ahhh, yep. that was it. thanks man

hartnn (hartnn):

sorry, typo, \(2\int 3^{-u}du = 2[\dfrac{3^{-u}}{(-1)\log 3}]+c\)

OpenStudy (anonymous):

where is the (-1) coming from?

hartnn (hartnn):

in derivatives, we multiply by the co-efficient of variable, similarly, in integration, we divide by the co-efficient of variable. so, here its -u thats why we need to divide by -1

OpenStudy (anonymous):

ahhh, ok. good explanation, thanks.

hartnn (hartnn):

welcome ^_^